# Oscillation-Reduced MXFP4 Training for Vision Transformers

### v1

@@ Abstract
In this work, we propose a novel training method TetraJet for a more accurate FP4 training. We comprehensively evaluate all of the quantizers involved in the training, and identify the weight oscillation problem in the forward pass as the main source of the degradation in MXFP4 training. Therefore, we introduce two novel methods, EMA Quantizer (Q-EMA) and Adaptive Ramping Optimizer (Q-Ramping), to resolve the oscillation problem. Extensive experiments on Vision Transformers demonstrate that TetraJet consistently outperforms the existing 4-bit training methods, and Q-EMA & Q-Ramping can provide additional enhancement by effectively reducing oscillation. We decreased the accuracy degradation by more than $50\%$ compared to the baseline, and can even achieve competitive performance compared to full precision training. The codes are available at https://github.com/thu-ml/TetraJet-MXFP4Training
@@ 1 Introduction
Low-precision training has emerged as a promising technique for accelerating the training process of large-scale neural networks. By quantizing tensors in both the forward and backward passes to lower-precision formats, low-precision training leverages specialized compute units in modern hardware to enhance computational efficiency significantly. While BF16 and FP16 precision remain the most widely used formats for deep learning training (Narang et al., 2017; Kalamkar et al., 2019), FP8 training (Sun et al., 2019; Micikevicius et al., 2022; NVIDIA, 2024c; Xi et al., 2024a) is becoming increasingly mature in these years, with successful application in training state-of-the-art large language models (Liu et al., 2024).
@@ 2 Related Work
Low-precision training has become a prominent technique in modern deep learning to speed up the training process. FP16 and BF16 (half-precision) training (Narang et al., 2017; Kalamkar et al., 2019) is currently the most common low-precision method. FP8 and INT8 training (Sun et al., 2019; Zhu et al., 2020; Micikevicius et al., 2022; Wortsman et al., 2023; NVIDIA, 2024c; Peng et al., 2023; Xi et al., 2024b, a; Liu et al., 2024) further improves efficiency and uses more fine-grained per-tensor / per-row / per-block quantization. When it comes down to 4-bit training (Sun et al., 2020; Chmiel et al., 2021; Xi et al., 2023), more techniques are being applied (e.g. Hadamard transformation) to compromise the degradation caused by the low representation ability. Still, their accuracy degradation is not negligible.
@@ 3.1 Preliminary
$X_{i}=\mathrm{round_{FP4}}\left(\frac{X_{i}}{S}\right),\quad X_{i}\approx P_{i}\cdot S.$ (1)
@@ 3.2 Quantization with Truncation-Free Scaling
Stochastic rounding is unbiased: $\mathbb{E}[\mathrm{round}_{S}(X_{i}/S)]=X_{i}/S$. We show the superiority of stochastic rounding in the ablation study in Sec. 7.3,
@@ 3.3 TetraJet Linear Layer
As a fine-grained format, doing MM with MXFP4 is more subtle than other coarser-grained formats such as per-tensor quantization. For hardware-accelerated MM to be possible, MXFP4 format requires quantization group shape to be $1\times 32$ for the first matrix and $32\times 1$ for the second matrix. Therefore, quantizer (1)(3)(5) should use $1\times 32$ group shape, and quantizer (2)(4)(6) should use $32\times 1$ group shape. This means that weight $\bm{\mathbf{W}}$, activation $\bm{\mathbf{X}}$, and gradient $\nabla_{\bm{\mathbf{Y}}}\mathcal{L}$ should be quantized along different axes in different quantizers. For example, the quantization block size of activation $\bm{\mathbf{X}}$ should be $1$ token $\times$ 32 channels in forward and $32$ tokens $\times$ 1 channel in backward.
@@ 4.2 Quantization Confidence of Weight Distribution
We can also define quantization confidence for a matrix $\bm{\mathbf{W}}$ as the average $\mathrm{QuantConf}(w)$ of all the element $w$ in $\bm{\mathbf{W}}$. The rationale behind this metric is that the closer a latent weight is to a quantization decision threshold, the more likely it is to oscillate, making it harder for the weight to converge to a stable FP4 value.
@@ 7.1 Vision Transformer Pre-Training
We do pre-training for DeiT-Tiny, DeiT-Small, and DeiT-Base (Touvron et al., 2020) using Facebook’s training recipe [^2] https://github.com/facebookresearch/deit, and pre-train Swin-Tiny and Swin-Small (Liu et al., 2021) based on the official implementation [^3] https://github.com/microsoft/Swin-Transformer. All the models are trained for 90 epochs on ImageNet1K (Russakovsky et al., 2015) with default training recipes. For Q-EMA & Q-Ramping, we show the insensitivity to their hyperparameter choice in Section C.3.
@@ 8 Conclusion
Appendix A Statistics to Measure Oscillation and Instability in MXFP4 Training
@@ A.1 Oscillation Ratio
Compare Oscillation Ratio and Previous Metric
@@ A.2 Quantization Confidence
We can also define quantization confidence for a matrix $\bm{\mathbf{W}}$ as the average $\mathrm{QuantConf}(w)$ of all the element $w$ in $\bm{\mathbf{W}}$.

### v2

@@ Abstract
In this work, we propose a novel training method TetraJet for a more accurate FP4 training. We comprehensively evaluate all of the quantizers involved in the training, and identify the weight oscillation problem in the forward pass as the main source of the degradation in MXFP4 training. Therefore, we introduce two novel methods, EMA Quantizer (Q-EMA) and Adaptive Ramping Optimizer (Q-Ramping), to resolve the oscillation problem. Extensive experiments on Vision Transformers demonstrate that TetraJet consistently outperforms the existing 4-bit training methods, and Q-EMA & Q-Ramping can provide additional enhancement by effectively reducing oscillation. We decreased the accuracy degradation by more than $50\%$ compared to the baseline, and can even achieve competitive performance compared to full precision training. The codes are available at https://github.com/thu-ml/TetraJet-MXFP4Training.
@@ 1 Introduction
Low-precision training has emerged as a promising technique for accelerating the training process of large-scale neural networks. By quantizing tensors in both the forward and backward passes to lower-precision formats, low-precision training leverages specialized compute units in modern hardware to enhance computational efficiency significantly. While BF16 and FP16 precision remain the most widely used formats for deep learning training (Narang et al., 2017; Kalamkar et al., 2019), FP8 training (Sun et al., 2019; Micikevicius et al., 2022; NVIDIA, 2024c; Xi et al., 2025) is becoming increasingly mature in these years, with successful application in training state-of-the-art large language models (Liu et al., 2024).
@@ 2 Related Work
Low-precision training has become a prominent technique in modern deep learning to speed up the training process. FP16 and BF16 (half-precision) training (Narang et al., 2017; Kalamkar et al., 2019) is currently the most common low-precision method. FP8 and INT8 training (Sun et al., 2019; Zhu et al., 2020; Micikevicius et al., 2022; Wortsman et al., 2023; NVIDIA, 2024c; Peng et al., 2023; Xi et al., 2024, 2025; Liu et al., 2024) further improves efficiency and uses more fine-grained per-tensor / per-row / per-block quantization. When it comes down to 4-bit training (Sun et al., 2020; Chmiel et al., 2021; Xi et al., 2023), more techniques are being applied (e.g. Hadamard transformation) to compromise the degradation caused by the low representation ability. Still, their accuracy degradation is not negligible.
@@ 3.1 Preliminary
$P_{i}=\mathrm{round_{FP4}}\left(\frac{X_{i}}{S}\right),\quad X_{i}\approx P_{i}\cdot S.$ (1)
@@ 3.2 Quantization with Truncation-Free Scaling
Stochastic rounding is unbiased: $\mathbb{E}[\mathrm{round}_{S}(X_{i}/S)]=X_{i}/S$. We show the superiority of stochastic rounding in the ablation study in Sec. 7.3.
@@ 3.3 TetraJet Linear Layer
As a fine-grained format, doing MM with MXFP4 is more subtle than other coarser-grained formats such as per-tensor quantization. For hardware-accelerated MM to be possible, MXFP4 requires quantization group shape to be $1\times 32$ for the first matrix and $32\times 1$ for the second matrix. Therefore, quantizers $Q^{(1)},Q^{(3)},Q^{(5)}$ should use $1\times 32$ group shape, and quantizers $Q^{(2)},Q^{(4)},Q^{(6)}$ should use $32\times 1$ group shape. This means that weight $\bm{\mathbf{W}}$, activation $\bm{\mathbf{X}}$, and gradient $\nabla_{\bm{\mathbf{Y}}}\mathcal{L}$ should be quantized along different axes in different quantizers. For example, the quantization block size of $\bm{\mathbf{X}}$ should be $1$ token $\times$ 32 channels in forward and $32$ tokens $\times$ 1 channel in backward.
@@ 4.2 Quantization Confidence of Weight Distribution
The rationale behind this metric is that the closer a latent weight is to a quantization decision threshold, the more likely it is to oscillate, making it harder for the weight to converge to a stable FP4 value.
@@ 7.1 Vision Transformer Pre-Training
We do pre-training for DeiT-Tiny, DeiT-Small, and DeiT-Base (Touvron et al., 2021) using Facebook’s training recipe [^2] https://github.com/facebookresearch/deit, and pre-train Swin-Tiny and Swin-Small (Liu et al., 2021) based on the official implementation [^3] https://github.com/microsoft/Swin-Transformer. All the models are trained for 90 epochs on ImageNet1K (Russakovsky et al., 2015) with default training recipes. For Q-EMA & Q-Ramping, we show the insensitivity to their hyperparameter choice in Appendix C.3.
@@ 8 Conclusion
Acknowledgment
The authors gratefully thank Pengle Zhang for the insightful discussions. This work was supported by the NSFC Project (No. 62376131), Beijing Natural Science Foundation (QY24258), Tsinghua University Initiative Scientific Research Program (Student Academic Research Advancement Program), and the High Performance Computing Center, Tsinghua University. J.Z is also supported by the XPlorer Prize.
@@ Acknowledgment
Appendix A Statistics for Measuring Oscillation and Training Instability
@@ A.1 Oscillation Ratio
Comparison of Oscillation Ratio and Previous Metric